OpenStudy (anonymous):

Determine the function f(x) satisfying the given conditions. f''(x) = 0 f'(5) = 1 f(4) = -2

OpenStudy (anonymous):

they want me to find and f'(x) and f(x)

OpenStudy (anonymous):

f''(x)=0 means your function is not going to be of higher grade than 1. From this you can assume that your function can be written as f(x)=ax+b f'(x)=a

OpenStudy (anonymous):

so it would be f(x)= x+5 and f'(x)=0

OpenStudy (anonymous):

right?

OpenStudy (anonymous):

If you df(x)/dx you have to remember the rules of derivatives.

OpenStudy (radar):

f'(x) when f(x) = x+5 would be 1 not zero

OpenStudy (anonymous):

f'(x)=x^0 --> 1

OpenStudy (anonymous):

am i supposed to find the antiderivative?

OpenStudy (radar):

Is this 3 problems, or are you saying one function to meet the three different conditions aobve??

OpenStudy (anonymous):

You have the function f(x), to get f'(x) you just derive f(x).

OpenStudy (anonymous):

just 1 problem radar

OpenStudy (radar):

f'(x) when x^0 is not 1, x^0=1 f' x^0=0 as it is a constant.

OpenStudy (anonymous):

so the function is f(x) is x-2?

OpenStudy (anonymous):

so the deriv is 1

OpenStudy (radar):

Is the first one asking for f'' like f double prime, derivative of the derivative?

OpenStudy (anonymous):

no just the first deriv

OpenStudy (radar):

Well the derivative of any constant is 0 as there is no "change" Let me think how the next two conditions play into this!?

OpenStudy (anonymous):

this is the format they want for f(x) = ax^2+bx+c

OpenStudy (anonymous):

Really?

OpenStudy (anonymous):

Because that does not go together with (f''(x) = 0).

OpenStudy (anonymous):

i just tried a previous problem: Determine the function f(x) satisfying the given conditions. Hint: To answer this question and the next one, think in terms of doing two successive antiderivatives. f''(x) = 16 f'(0) = 7 f(0) = 4 answer: f(X)=8x^2+7x+4

OpenStudy (anonymous):

If you got a different question please post a new question.

OpenStudy (anonymous):

no this just an example

OpenStudy (radar):

Try this: f(x)= (1/2)( x^2) - (1/4)( x^2)

OpenStudy (anonymous):

You could use the form f(x)=ax^2+bx+c, that just means a has to be 0 in your example.

OpenStudy (anonymous):

and you'll get f(x)=bx+c which is basically the same thing as f(x)=ax+b

OpenStudy (radar):

f' would be !!! never mind, back to the drawing board lol

OpenStudy (anonymous):

i think f'(x) has to be a constant. so f'(x)=C for some value of c.

OpenStudy (anonymous):

f''(x)=0 --> f(x)=bx+c --> f'(x)=b f'(5)=1 --> f'(5)=1 --> b=1 f(4)=-2 --> f(4)=4+c=-2 --> c=-6 f'(x)=1 f(x)=x-6

OpenStudy (anonymous):

yes thats right

OpenStudy (radar):

From what you listed as a previous example. The first one is a double derivative f" The second one is a single derivative The third one is the function itself. Now look at it that way.

OpenStudy (anonymous):

ok so u have to find 2nd deriv then work your way down to the function

OpenStudy (anonymous):

its all in reverse! lol

OpenStudy (radar):

That is what it appears to be, and the function 8x^2+7x+4 convinces me.

OpenStudy (anonymous):

ok thanx

OpenStudy (radar):

Hopefully your text will show you the procedure to do that.....It does not seem to be a trivial task.